# Common or specific source, features or scores; it is all a matter of information

### v1

@@ 2.1 Two hypotheses and two actions
We consider two mutually exclusive hypotheses $H_{1}$ and $H_{2}$ with ${\mathbb{P}}(H_{1})+{\mathbb{P}}(H_{2})=1$, and actions $A_{1}$ and $A_{2}$. We assume that $c_{11}=c_{22}=0$, so that $A_{i}$ is the ‘correct’ action if $H_{i}$ is true and can be executed without costs. Since Bayes decisions will remain the same when all costs are multiplied by the same factor, we can without loss of generality assume that let $c_{21}=1$ and $c_{12}=\gamma>0$. The expected cost of $A_{1}$ is then equal to $\gamma{\mathbb{P}}(H_{2})$, and the expected cost of $A_{2}$ is equal to ${\mathbb{P}}(H_{1})$. Hence, if we were to decide at this point, one chooses $A_{1}$ if and only if
@@
that is, if the prior odds $r$ are at least $\gamma$. We are now in the same setting as Section 3 of [28] with the constant $\gamma$ playing the role of the threshold value $Th$.
@@
for this probability under $H_{1}$. Furthermore, it is known that (cf. the proof of Proposition 2.4.2 in [18])
@@
If $H_{1}$ is true, the probability that the posterior odds exceed $\gamma$ is $s$. If that happens, we choose $A_{1}$ with no costs. If the posterior odds do not exceed $\gamma$, then we choose $A_{2}$ with costs 1, something which happens with probability $1-s$. Hence, under $H_{1}$ the expected costs are $1-s$.
If $H_{2}$ is true, we only make costs if the posterior odds exceed $c$, and this happens with probability at most $s\cdot r/\gamma$. It follows that the expected posterior costs of the BD are at most
@@
Since the prior expected costs are ${\mathbb{P}}(H_{1})$, we see indeed that the expected costs of the BD have decreased. A similar reasoning holds when $r\geq\gamma$. This proves the result.
@@ 2.2 The general case
Recall that our claim is that
${\mathbb{E}}(c(\pi(E))\leq c(\pi),$ (2.6)
and we now prove this.
We start with a general observation. Let, for $k=1,2,\ldots$, $p_{k}=(p_{k,1},\ldots,p_{k,n})$ be probability vectors, and let $t_{1},t_{2},\ldots$ be non-negative numbers such that $\sum_{k=1}^{\infty}t_{k}=1$. Then $\sum_{k}^{\infty}t_{k}p_{k}$ is again a probability vector, and we have
$\begin{split}c\left(\sum_{k=1}^{\infty}t_{k}p_{k}\right)&=\min_{i}\sum_{j=1}^{n}c_{ij}\sum_{k=1}^{\infty}t_{k}p_{k,j}\\ &=\min_{i}\sum_{k=1}^{\infty}\sum_{j=1}^{n}c_{ij}t_{k}p_{k,j}\\ &=\min_{i}\sum_{k=1}^{\infty}t_{k}\sum_{j=1}^{n}c_{ij}p_{k,j}\\ &\geq\sum_{k=1}^{\infty}t_{k}\min_{i}\sum_{j=1}^{n}c_{ij}p_{k,j}\\ &=\sum_{k=1}^{\infty}t_{k}c(p_{k}).\\ \end{split}$
@@
It remains to show the, intuitively obvious, fact that ${\mathbb{E}}[\pi(E)]=\pi$, i.e., that the expectation of the posterior probability distribution $\pi(e)$ is equal to the prior $\pi$. For this, we use that the prior and posterior are related via the likelihood ratio; this part of the argument replaces (2.5) in Section 2.1.
Writing $H_{r}$ for the union of $H_{2},\ldots,H_{n}$, $(\pi_{1},\pi_{r})$ for the priors, $LR$ for the likelihood ratio of $H_{1}$ relative to $H_{r}$, and $\ell_{i}(\cdot):={\mathbb{P}}(LR=x\mid H_{i})$, we have that the expected posterior probability of $H_{1}$ is given by
$\sum_{e}p(e)\frac{LR(e)\pi_{1}}{LR(e)\pi_{1}+\pi_{r}}$ $=$ $\sum_{x}{\mathbb{P}}(LR=x)\frac{x\pi_{1}}{x\pi_{1}+\pi_{r}}$
$=$ $\sum_{x}\left(\pi_{1}{\mathbb{P}}(LR=x\mid H_{1})+\pi_{r}{\mathbb{P}}(LR=x\mid H_{r})\right)\frac{x\pi_{1}}{x\pi_{1}+\pi_{r}}$
$=$ $\sum_{x}\left(\pi_{1}\ell_{1}(x)+\pi_{r}\ell_{r}(x)\right)\frac{x\pi_{1}}{x\pi_{1}+\pi_{r}}$
$=$ $\sum_{x}\left(\pi_{1}x\ell_{r}(x)+\pi_{r}\ell_{r}(x)\right)\frac{x\pi_{1}}{x\pi_{1}+\pi_{r}}$
$=$ $\sum_{x}\ell_{r}(x)\pi_{1}x$
$=$ $\sum_{x}\ell_{1}(x)\pi_{1}=\pi_{1}.$
This proves the claim.
This information-theoretical perspective is the key for understanding the distinction between the common source versus the specific source scenario on the one hand, and the difference between score-based and feature-based on the other. In both comparisons, the issue is that one of the alternatives is based on less information than the other. Inequality (2.2) expresses that on average, the better-informed situations leads to better decisions. That does not mean that the less informed LRs are ‘wrong’ in whatever sense, only that they are based on less information. If this, however, is all the information that is available, then using this information is the right thing to do.
So far, these assessments have all been qualitative. Full features are better than scores, but to what extent? Or put differently, how bad is it to use less information? Measuring performance is often a very contextual matter, but in general it is well known [18] that, when comparing hypotheses $H_{1}$ and $H_{2}$,
@@
Now, suppose that we consider two hypotheses $H_{1}$ and $H_{2}$ with some prior $\pi$ and data $e$, and that we have obtained $\pi(g(e))$ by computing $LR(g(e))$ for some $g(e)$. Then, we can update $\pi(g(e))$ to $\pi(e)$ by calculating $LR(e\mid g(e))$, so that the two inequalities above apply to the computation of the LR for $e\mid g(e)$. Hence, if $H_{1}$ were true, then the probability that $LR(e\mid g(e))$ is much smaller than $LR(g(e))$, is bounded and so is the probability that $LR(e\mid g(e))$ is much larger than $LR(g(e))$ if $H_{2}$ is true.
Hence, if for example $g(e)$ provides strong evidence for either hypothesis, then the probability that taking the full data $e$ into account will incorrectly point strongly in the other direction is bounded. We next illustrate these observations with an extensive example.
@@ 2.3 Example: DNA kinship LRs
In this example, the goal is to investigate whether two persons are siblings or unrelated. To that end, we consider the DNA profile of a person as a ‘trace measurement’ of the DNA of their parents. Thus, the population sources is the population of pairs of man and women (assumed to be always unrelated to each other), and sources give rise to measurements in the form of a DNA profile of a child of theirs. This model is well understood: sources are described by a population genetically obtained probability distribution on the DNA profiles of persons in the general population, and children are obtained by Mendelian inheritance (perhaps enriched with a mutation model).
Now, suppose we have two individuals $X$ and $Y$ that are alleged siblings. Write $H_{0}$ for the hypothesis that they are full siblings, and $H_{1}$ for the hypothesis that they are unrelated. There are several types of measurements we could have: $e_{x}$ (the DNA profile of $X$), $e_{y}$ (the DNA profile of $Y$), and, say, $\theta_{x}$ (the DNA profiles of the parents of $X$).
@@
But this is absolutely no reason to dismiss, or in any way distrust, the LR based on the profiles of the alleged siblings only. When only $(e_{x},e_{y})$ are available, the evidence for being siblings versus unrelated is given by $LR(e_{x},e_{y})$. Note also that when this LR is large, the posterior probability for being siblings increases as well, and if $X$ and $Y$ are indeed siblings then the scenario that we just sketched (where additionally obtaining the profiles of the parents of $X$ and excluding the relationship) is not possible.
@@
Now, instead of processing the profiles $(e_{x},e_{y})$ we could consider a score $g(e_{x},e_{y})$, for example counting the number of alleles that the profiles have in common. Again, if only this score were known to us, the evidence is then $LR(g(e_{x},e_{y}))$. There is nothing inherently wrong with this LR, but since the pair $(e_{x},e_{y})$ carries more information, we would rather assign a LR based on the full data, then only on the score.
@@
To illustrate this, we have run a simulation experiment, generating all data just described: both for full siblings and unrelated individuals, we count the number of shared alleles between $e_{x}$ and $e_{y}$, we compute the likelihood ratio $LR(e_{x},e_{y})$ comparing being siblings to being unrelated, and we compute the LR for $e_{x}$ to be a child of the parents of $X$, versus unrelated to both of them. We have done so based on Dutch allele frequencies ([30]) and using various DNA multiplexes: one with 10, and one with 15 autosomal loci. In all cases the profiles were simulated based on 15 loci and then calculations were done for these, as well for the subset of 10 loci. We simulated 10,000 cases. Note that these are, nowadays, rather small numbers of loci; we chose them purely for illustration purposes.
@@
First of all, suppose that we work on 10 loci. We obtain likelihood ratios $LR(g(e_{x},e_{y}))$ for being siblings, versus unrelated, based on the empirical number $g(e_{x},e_{y})$ of alleles shared on 10 loci. We can compare these to the LRs obtained when $LR(e_{x},e_{y})$ is computed, and the result is displayed in Figures 2.2 and 2.2.
Figure 2.1: $\log_{10}(LR(e_{x},e_{y}))$ versus $\log_{10}(LR(g(e_{x},e_{y})))$ for actual siblings.
Figure 2.2: $\log_{10}(LR(e_{x},e_{y}))$ versus $\log_{10}(LR(g(e_{x},e_{y})))$ for unrelated.
In these figures, the vertical lines correspond to a number of shared alleles, with neutral evidence obtained for 9 (out of possibly 20) shared alleles. We see in these figures that, when the profiles are evaluated, the LRs of course change, and have a tendency to support the correct hypothesis.
@@
Indeed, comparing $LR(e_{x},e_{y})$ with $LR(g(e_{x},e_{y}))$ by considering their quotient, we see that $LR(e_{x},e_{y})/LR(g(e_{x},e_{y}))$ respects the bounds (2.8) and (2.9). In Figures 2.4 and 2.4 we plot the distribution of the difference between the two LRs on a logarithmic scale. Clearly, both for siblings and unrelated individuals the evidence tends to more strongly towards the hypothesis that is actually true. There are, also, cases where this does not happen, which is inevitable and to be expected. The frequency with which this occurs, however, is bounded by (2.8) and (2.9).
Figure 2.3: $\log_{10}(LR(e_{x},e_{y}))-\log(LR(g(e_{x},e_{y})))$ for actual siblings.
Figure 2.4: $\log_{10}(LR(e_{x},e_{y}))-\log_{10}(LR(g(e_{x},e_{y})))$ for unrelated.
@@
Thus, we see that when we process the profiles $(e_{x},e_{y})$ instead of the number of shared alleles $g(e_{x},e_{y})$, the LRs change. That is no reason to say that, in hindsight, the LR based on $g(e_{x},e_{y})$ is incorrect; it is only a sub-optimal way to treat the data. But so is $(e_{x},e_{y})$ when more loci become available: comparing the LRs for $(e_{x},e_{y})$ on 15 with those restricted on 10 loci gives similar changes, as shown in Figures 2.6 and 2.6.
Figure 2.5: Difference in $\log_{10}(LR(e_{x},e_{y}))$ on 15 versus 10 loci, for actual siblings.
Figure 2.6: Difference in $\log_{10}(LR(e_{x},e_{y}))$ on 15 versus 10 loci for unrelated.
@@
The most dramatic changes in LR occur, of course, if the parents of $X$ become known and the question of whether $X$ and $Y$ are siblings reduces to whether $Y$ is a child of the parents of $X$. In that case, going back to 10 loci, we get the changes in LR displayed in Figures 2.8 and 2.8.
Figure 2.7: Difference in $\log_{10}(LR)$ based on profiles, with or without parents of $X$, for actual siblings.
Figure 2.8: Difference in $\log_{10}(LR)$ based on profiles, with or without parents of $X$ for unrelated.
@@
In conclusion, we see in this example how the different “LR systems” can all be defined in a context where they are not traditionally thought of as such; indeed, all these LRs simply differ in which information is available, and whether the full information is processed or only a function of it. Of course, when such a calculation is done, it is of utmost importance that the forensic report contains a full description of the considered hypotheses and the processed data. A LR based on the profiles of the alleged siblings alone, ignoring the profiles of the parents, is not wrong: it expresses the information obtained from what is taken into account. At the same time, of course, if more data is available for which a statistical model is available, any analysis that ignores these data is sub-optimal.

### v2

@@ 2.1 Two hypotheses and two actions
We consider two mutually exclusive hypotheses $H_{1}$ and $H_{2}$ with ${\mathbb{P}}(H_{1})+{\mathbb{P}}(H_{2})=1$, and actions $A_{1}$ and $A_{2}$. We assume that $c_{11}=c_{22}=0$, so that $A_{i}$ is the ‘correct’ action if $H_{i}$ is true and can be executed without costs. Since Bayes decisions will remain the same when all costs are multiplied by the same factor, we can without loss of generality assume that $c_{21}=1$ and $c_{12}=\gamma>0$. The expected cost of $A_{1}$ is then equal to $\gamma{\mathbb{P}}(H_{2})$, and the expected cost of $A_{2}$ is equal to ${\mathbb{P}}(H_{1})$. Hence, if we were to decide at this point, one chooses $A_{1}$ if and only if
@@
that is, if the prior odds $r$ are at least $\gamma$. We are now in the same setting as Section 3 of vergeer with the constant $\gamma$ playing the role of the threshold value $Th$.
@@
for this probability under $H_{1}$. Furthermore, it is known that (cf. the proof of Proposition 2.4.2 in meesterslootenbook)
@@
If $H_{1}$ is true, the probability that the posterior odds exceed $\gamma$ is $s$. If that happens, we choose $A_{1}$ with no costs. If the posterior odds do not exceed $\gamma$, then we choose $A_{2}$ with costs 1, something which happens with probability 1s1-s. Hence, under $H_{1}$ the expected costs are 1s1-s.
If $H_{2}$ is true, we only make costs if the posterior odds exceed $\gamma$, and this happens with probability at most $s\cdot r/\gamma$. It follows that the expected posterior costs of the BD are at most
@@
Since the prior expected costs are ${\mathbb{P}}(H_{1})$, we see indeed that the expected costs of the BD are at most the same as based on the prior probabilities. A similar reasoning holds when $r\geq\gamma$. This proves Theorem 2.1 for this case.
@@ 2.2 The general case
To prove Theorem 2.1 we start with a general observation. Let, for $k=1,2,\ldots$, $p_{k}=(p_{k,1},\ldots,p_{k,n})$ be probability vectors, and let $t_{1},t_{2},\ldots$ be non-negative numbers such that $\sum_{k=1}^{\infty}t_{k}=1$. Then $\sum_{k}^{\infty}t_{k}p_{k}$ is again a probability vector, and we have
[…]
@@
Alternatively, one can apply Jensen’s inequality to arrive at the same conclusion.
It remains to show that ${\mathbb{E}}[\pi(E)]=\pi$, i.e., that the expectation of the posterior probability distribution $\pi(e)$ is equal to the prior $\pi$. But this follows immediately from the fact that
${\mathbb{E}}({\mathbb{P}}(H_{i}\mid E))=\sum_{e}{\mathbb{P}}(H_{i}\mid e){\mathbb{P}}(e)={\mathbb{P}}(H_{i}),$
for all $i=1,\ldots,n$. This proves Theorem 2.1.
This information-theoretical perspective is the key for understanding the distinction between the common source versus the specific source scenario on the one hand, and the difference between score-based and feature-based on the other. In both comparisons, the issue is that one of the alternatives is based on less information than the other. Theorem 2.1 expresses that on average, the better-informed situations leads to better decisions. That does not mean that the less informed LRs are ‘wrong’ in whatever sense, only that they are based on less information. If this, however, is all the information that is available, then using this information is the right thing to do.
So far, these assessments have all been qualitative. Full features are better than scores, but to what extent? Or put differently, how bad is it to use less information? Measuring performance is often a very contextual matter, but in general it is well known meesterslootenbook that, when comparing hypotheses $H_{1}$ and $H_{2}$,
@@
Now, suppose that we consider two hypotheses $H_{1}$ and $H_{2}$ with some prior $\pi$ and data $e$, and that we have obtained $\pi(g(e))$ by computing $LR(g(e))$ for some $g(e)$. Then, we can update $\pi(g(e))$ to $\pi(e)$ by calculating $LR(e\mid g(e))$, so that the two inequalities above apply to the computation of the LR for $e\mid g(e)$. Hence, if $H_{1}$ were true, then the probability that $LR(e\mid g(e))$ is more than a factor $t$ smaller than $LR(g(e))$, is bounded by $1/t$, and so is the probability that $LR(e\mid g(e))$ is more than a factor $t$ larger than $LR(g(e))$ if $H_{2}$ is true.
Hence, if for example $g(e)$ provides strong evidence for either hypothesis, then the probability that taking the full data $e$ into account will incorrectly (or rather, unfortunately) point strongly in the other direction is bounded as just described. We next illustrate these observations with an extensive example.
@@ 2.3 Example: DNA kinship LRs
In this example, the goal is to investigate whether two persons are siblings or unrelated. To that end, we consider the DNA profile of a person as a ‘trace measurement’ of the DNA of their parents. Thus, the population of sources is the population of pairs of man and women (assumed to be always unrelated to each other), and sources give rise to measurements in the form of a DNA profile of a child of theirs. This model is well understood: DNA-profiles of parents are described by a probability distribution on the DNA profiles of persons in the general population, and children are obtained by Mendelian inheritance (perhaps enriched with a mutation model).
Now, suppose we have two individuals $X$ and $Y$ who are either siblings or unrelated. Write $H_{0}$ for the hypothesis that they are full siblings, and $H_{1}$ for the hypothesis that they are unrelated. Since we assume that these two hypotheses are exhaustive (i.e., have total probability equal to one), this example directly connects to Section 2.1. Furthermore, we assume the following types of measurements: $e_{x}$ (the DNA profile of $X$), $e_{y}$ (the DNA profile of $Y$), and, say, $\theta_{x}$ (the DNA profiles of the parents of $X$).
@@
But this is absolutely no reason to dismiss, or in any way distrust, the LR based on the profiles of the alleged siblings only. When only $(e_{x},e_{y})$ are available, the evidence for being siblings versus unrelated is given by $LR(e_{x},e_{y})$. Note also that when this LR becomes larger, the posterior probability for being siblings increases as well, and if $X$ and $Y$ are indeed siblings then the scenario that we just sketched (additionally obtaining the profiles of the parents of $X$ and excluding the relationship with $Y$) is not possible.
@@
Now, instead of processing the profiles $(e_{x},e_{y})$ we could consider a score $g(e_{x},e_{y})$, for example counting the number of alleles that the profiles have in common. Again, if only this score were known to us, the evidence is then $LR(g(e_{x},e_{y}))$. There is nothing inherently wrong with this LR, but since the pair $(e_{x},e_{y})$ carries more information, we would rather assign a LR based on the full data, than only on the score.
@@
To illustrate this, we have run a simulation experiment, generating all data just described: both for full siblings and unrelated individuals, we count the number $g(e_{x},e_{y})$ of shared alleles between $e_{x}$ and $e_{y}$, we compute the likelihood ratio $LR(e_{x},e_{y})$ comparing being siblings to being unrelated, and we compute the LR for $e_{y}$ to be a child of the parents of $X$, versus unrelated to both of them. We have done so based on Dutch allele frequencies (FLDOfreqs) and using various DNA multiplexes: one with 10, and one with 15 autosomal loci. In all cases the profiles were simulated based on 15 loci and then calculations were done for these, as well for the subset of 10 loci. We simulated 500,000 cases. Note that these are, nowadays, rather small numbers of loci; we chose them purely for illustration purposes.
@@
First of all, suppose that we work on 10 loci. We obtain likelihood ratios $LR(g(e_{x},e_{y}))$ for being siblings, versus unrelated, based on the empirical number $g(e_{x},e_{y})$ of alleles shared on 10 loci. We can compare these to the LRs obtained when $LR(e_{x},e_{y})$ is computed, and the result is displayed in Figures 2.1.
(a) true relationship: siblings
(b) true relationship: unrelated
Figure 2.1: $\log_{10}(LR(e_{x},e_{y}))$ (based on DNA profiles) versus $\log_{10}(LR(g(e_{x},e_{y})))$ (based on number of shared alleles). Each violin plot represents the $LR(e_{x},e_{y})$ for profiles whose number of shared alleles is displayed in the plot.
In these figures, we have placed violin plots representing the distribution of $LR(e_{x},e_{y})$ placed at the values $LR(g(e_{x},e_{y}))$ corresponding to a number $g(e_{x},e_{y})$ of shared alleles, and we see for example that neutral evidence is obtained for 9 (out of possibly 20) shared alleles. We see in these figures that, when the profiles are evaluated, the $LR(e_{x},e_{y})$ are of course different from $LR(g(e_{x},e_{y}))$, and have a tendency to more strongly support the correct hypothesis. It may, of course, happen that the LR based on $(e_{x},e_{y})$ supports the other hypothesis compared to the LR based on $g(e_{x},e_{y})$. But we also see that this becomes less likely for larger $|Log_{10}LR(g(e_{x},e_{y})|$, as predicted. Finally we note that each violin plot has the same width irrespective of the number of instances of $g(e_{x},e_{y})$.
@@
In Figure 2.2 we plot the distribution of the difference between the two LRs on a logarithmic scale. Indeed, comparing $LR(e_{x},e_{y})$ with $LR(g(e_{x},e_{y}))$ by considering their quotient, we see that $LR(e_{x},e_{y})/LR(g(e_{x},e_{y}))$ respects the bounds (2.5) and (2.6). Clearly, both for siblings and unrelated individuals the evidence tends more strongly towards the hypothesis that is actually true. There are, also, cases where this does not happen, which is inevitable and to be expected. The frequency with which this occurs, however, is bounded by (2.5) and (2.6).
(a) Actual siblings
(b) Unrelated pairs
Figure 2.2: Cumulative distribution function of $\log_{10}(LR(e_{x},e_{y}))-\log_{10}(LR(g(e_{x},e_{y})))$ (black), and theoretical bound (dashed).
@@
Thus, we see that when we process the profiles $(e_{x},e_{y})$ instead of the number of shared alleles $g(e_{x},e_{y})$, the LRs change. That is no reason to say that, in hindsight, the LR based on $g(e_{x},e_{y})$ is incorrect; it is only a sub-optimal way to treat the data. But so is $(e_{x},e_{y})$ when more loci become available: comparing the LRs for $(e_{x},e_{y})$ on 15 with those restricted on 10 loci gives similar changes, as shown in Figures 3(a) and 3(b). Note that, since for these loci the genetic data of unrelated persons are independent on the loci considered, and so is the inheritance of alleles towards offspring, these graphs also represent the likelihood ratio distribution on the five additional loci.
(a) Actual siblings
(b) Unrelated pairs
Figure 2.3: Difference in $\log_{10}(LR(e_{x},e_{y}))$ on 15 versus 10 loci.
@@
The most dramatic changes in LR occur, of course, if the parents of $X$ become known and the question of whether $X$ and $Y$ are siblings reduces to whether $Y$ is a child of the parents of $X$. In that case, going back to 10 loci, we get the changes in LR displayed in Figure 2.4.
(a) Actual siblings.
(b) Unrelated pairs
Figure 2.4: Difference in $\log_{10}(LR)$ based on profiles, with or without parents of $X$.
@@
In conclusion, we see in this example how the different “LR systems” can all be defined in a context where they are not traditionally thought of as such; indeed, all these LRs simply differ in which information is available, and whether the full information is processed or only a function of it. They all compare the hypothesis “the source of $Y$ is $S_{x}$, the source of XX” to the alternative “the source of $Y$ is a random unknown source”, either on the basis of $g(e_{x},e_{y})$, $(e_{x},e_{y})$ or $(e_{x},e_{y},\theta_{x})$. Of course, when such a calculation is done, it is of utmost importance that the forensic report contains a full description of the considered hypotheses and the processed data. A LR based on the profiles of the alleged siblings alone, ignoring the profiles of the parents, is not wrong: it expresses the information obtained from what is taken into account. At the same time, of course, if more data is available for which a statistical model is available, any analysis that ignores these data is sub-optimal. Although we assumed only two options throughout this example, the general conclusions also hold in case of more or two non-exhaustive hypotheses. In those cases, the LRs calculated above only address the strength of evidence relative to two hypotheses at hand. The point remains that more information will, on average, never yield worse decisions.